OpenStudy (anonymous):

Solve the given nonlinear inequality. Write the solution set using interval notation. x2 − 3x + 4 x + 1 ≤ 1

OpenStudy (anonymous):

x-2/3x+4-6x/x+1? I am assuming the question is (x - 2)/(3x + 4) - 6x/(x + 1) (x - 2)/(3x + 4) - 6x/(x + 1) = 0 Mulitply both terms by (3x + 4) (X - 2)(3x + 4) /(3x + 4) - 6x(3x + 4)/(x + 1) = 0 Simplify (x - 2) - 6x(3x + 4)/(x + 1) = 0 Mulitply both terms by (x + 1) (x - 2)(X + 1) - 6x(x + 1)(3x + 4)/(x + 1) Simplify x - 2)(X + 1) - 6x(3x + 4) = 0 Muliply out x^2 - x - 2 - 18x^2 -24x = 0 Collect like terms -17x^2 -25x -2 = 0 Multiply by -1 17x^2 + 25x + 2 = 0